OpenStudy (anonymous):

y=1/x^2 find the concavity, where its increasing and decreasing, the point of inflection, and graph it.

jimthompson5910 (jim_thompson5910):

If \[\Large f(x) = \frac{1}{x^2}\] then what is \(\Large f'(x)\) equal to?

OpenStudy (anonymous):

what do you mean

jimthompson5910 (jim_thompson5910):

are you able to find the derivative?

OpenStudy (anonymous):

hehe sounds like a nope

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ \frac{ 1 }{ (x+h) ^{2} }-\frac{ 1 }{ x ^{2} } }{ h }\]

OpenStudy (solomonzelman):

~ find f''(x). ~ set f''(x)=0 and solve for x ~ this will give you the x. so find the points on the function of these X-s. { call inflections x=a b c then the points of inflections are (a,f(a)) , (b,f(b)) , (c,f(c)). }

OpenStudy (solomonzelman):

to differentiate, re-write it as x^-2, and use the power rule.

OpenStudy (anonymous):

to find increasing and decreasing use f ' (x)=0 to find concavity and inflection point use f '' (x)=0

OpenStudy (anonymous):

Hope the attached plot from Mathematica 9 will help.

OpenStudy (solomonzelman):

lol, failed to load... :P